### v2

@@ 4 Convergence Analysis of The P-BCD Algorithm
Let $h:\mathbb{R}^{m\times n}\to\mathbb{R}$ be defined by $h(X)=\frac{1}{2}\|X-PQ^{\top}\|_{F}^{2}$, where $Q\in\mathbb{R}^{n\times m}$, $P\in\mathbb{R}^{m\times m}$ and $m\geq n$. If $X\in\mathbb{S}_{m,n}$ and
(17) $\bar{X}=\operatorname{Proj}_{\mathbb{S}_{m,n}}(X-\tau(X-PQ^{\top})),$
where $\tau\in(0,1)$, then we have $\bar{X}\in\mathbb{S}_{m,n}$ and
(18) $h(X)-h(\bar{X})\geq\frac{\tau^{-1}-1}{2(\tau^{-1}+1+\theta)^{2}}\|(\mathbf{I}_{m}-X{X}^{\top})\nabla h(X)\|_{F}^{2},$
where $\theta=\|PQ^{\top}\|_{2}$.
@@
Let $H$ be defined as in (6). Let $\{\mathcal{Z}^{k}\}$ be the sequence generated by Algorithm 1 and $[{X}_{i}^{k+1}]$ be computed by (9). Then $[X_{i}^{k+1}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and the following inequality holds:
(19) $\begin{split}&H([X_{1}^{k}],[X_{2}^{k}],[X_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([X_{1}^{k+1}],[X_{2}^{k+1}],[X_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}\end{split}$
@@
(20) $\left\|[H_{i}^{k}]^{\top}[{X}_{i}^{k}]-{[{X}_{i}^{k}]}^{\top}[H_{i}^{k}]\right\|_{F}\leq c_{4}\|[{X}_{i+1}^{k}]-[{X}_{i}^{k}]\|_{F},$
where $c_{3}=\frac{{\tilde{\alpha}_{X}}^{-1}-1}{2({\tilde{\alpha}_{X}}^{-1}+1+\theta_{\max})^{2}}>0$, $c_{4}=2(\alpha_{X}\sqrt{n_{\max}}+\theta_{\max})>0$, $\theta_{\max}=\max_{ijk}$ $\|[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{2}$, $n_{\max}=\max\{n_{1},n_{2},n_{3}\}$, and
$[H_{1}^{k}]$ $=\nabla_{[X_{1}]}H([{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$[H_{2}^{k}]$ $=\nabla_{[X_{2}]}H([{X}_{1}^{k+1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$[H_{3}^{k}]$ $=\nabla_{[X_{3}]}H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
@@
$[H_{i}^{k}]=([{X}_{i}^{k}][{Q}_{i}^{k}]-[{P}_{i}^{k}])[{Q}_{i}^{k}]^{\top},\quad i=1,2,3.$
@@
It follows from Lemma 4.5 that $[{X}_{i}^{k+1}]\in\mathbb{S}_{m_{i},n_{i}}$.
To show the first inequality holds, the update of $[{X}_{i}^{k+1}]$ in (9) can be viewed as the iteration in (17) with $h([X])=\frac{1}{2}\|[X]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\|_{F}^{2}$, $[X]=[{X}_{i}^{k}]$, $[\bar{X}]=[{X}_{i}^{k+1}]$ and $\tau=\tilde{\alpha}_{X}=\frac{\delta}{\delta+\alpha_{X}}$. Then for $i=1$ we have
$H([{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$-H([{X}_{1}^{k+1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$=$ $\frac{1}{2}\|[{X}_{i}^{k+1}][{Q}_{i}^{k}]-[{P}_{i}^{k}]\|_{F}^{2}-\frac{1}{2}\|[{X}_{i}^{k}][{Q}_{i}^{k}]-[{P}_{i}^{k}]\|_{F}^{2}$
$=$ $\frac{1}{2}\|[{X}_{i}^{k+1}]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\|_{F}^{2}-\frac{1}{2}\|[{X}_{i}^{k}]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\|_{F}^{2}$
$\geq$ $c_{ij}^{k}\left\|\left([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top}\right)\left([{X}_{i}^{k}]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)\right\|_{F}^{2},$
where $c_{ij}^{k}=\frac{{\tilde{\alpha}_{X}}^{-1}-1}{2({\tilde{\alpha}_{X}}^{-1}+1+\|[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{2})^{2}}\geq c_{3}>0$ and $\theta_{\max}$ is bounded, since the sequence $\{\mathcal{Z}^{k}\}$ is bounded. Using $[{X}_{i}^{k}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and $[{X}_{i}^{k+1}]\in[\mathbb{S}_{m_{i},n_{i}}]$, we can obtain the fourth line above and rewrite part of the last line as follows
$\left([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top}\right)\left([{X}_{i}^{k}]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$
$=$ $\left([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top}\right)\left([{X}_{i}^{k}][{Q}_{i}^{k}]-[{P}_{i}^{k}]\right)[{Q}_{i}^{k}]^{\top}$
$=$ $\left([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top}\right)[H_{i}^{k}].$
@@
$\begin{split}&H([{X}_{1}^{k}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\mathbf{I}_{m_{1}}]-[{X}_{1}^{k}]{[{X}_{1}^{k}]}^{\top}\right)[H_{1}^{k}]\right\|_{F}^{2}.\end{split}$
@@
$\begin{split}&H([{X}_{1}^{k+1}],[{X}_{2}^{k}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\mathbf{I}_{m_{2}}]-[{X}_{2}^{k}]{[{X}_{2}^{k}]}^{\top}\right)[H_{2}^{k}]\right\|_{F}^{2}\end{split}$
@@
$\begin{split}&H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([{X}_{1}^{k+1}],[{X}_{2}^{k+1}],[{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\mathbf{I}_{m_{3}}]-[{X}_{3}^{k}]{[{X}_{3}^{k}]}^{\top}\right)[H_{3}^{k}]\right\|_{F}^{2}.\end{split}$
@@
Next, we show the second inequality holds. According to the update of $[{X}_{i}^{k+1}]$ in (9), if we let
$[U^{k}][\Sigma^{k}][V^{k}]^{\top}=[{X}_{i}^{k}]-\tilde{\alpha}_{X}\left([{X}_{i}^{k}]-[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)=\frac{1}{\alpha_{X}+\delta}\left(\alpha_{X}[X_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right),$
then the update $[{X}_{i}^{k+1}]$ $=[U^{k}][V^{k}]^{\top}$. Hence, we have
(21) $[{X}_{i}^{k+1}]^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$ $=\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}[{X}_{i}^{k+1}]$
$=(\alpha_{X}+\delta)[V]^{k}[\Sigma^{k}][V^{k}]^{\top}.$
Then using $[{X}_{i}^{k}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and $[{X}_{i}^{k+1}]\in[\mathbb{S}_{m_{i},n_{i}}]$, we can rewrite
$[H_{i}^{k}]^{\top}[{X}_{i}^{k}]-{[{X}_{i}^{k}]}^{\top}[H_{i}^{k}]$
$=$ $\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}[{X}_{i}^{k}]-[{X}_{i}^{k}]^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$
$=$ $\left([{X}_{i}^{k+1}]-[{X}_{i}^{k}]\right)^{\top}\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)$
$-\left(\alpha_{X}[{X}_{i}^{k}]+\delta[{P}_{i}^{k}][{Q}_{i}^{k}]^{\top}\right)^{\top}\left([{X}_{i}^{k+1}]-[{X}_{i}^{k}]\right),$
where the last equation is obtained by (21). Taking the Frobenius norm of both sides, we obtain (20), where $2\|\alpha_{X}[{X}_{i}^{k}]^{(j)}+\delta[{P}_{i}^{k}]^{(j)}([{Q}_{i}^{k}]^{(j)})^{\top}\|_{F}\leq c_{4}$.
@@
${0}=\mathcal{L}^{\tilde{k}}+\mathcal{S}^{\tilde{k}}-\mathcal{D}+\delta\mathcal{W}_{\mathbf{R}}\odot\mathcal{L}^{\tilde{k}}-\delta\mathbf{R}^{\top}([\mathcal{Y}^{\tilde{k}}]),$
where $[\tilde{\mathcal{O}}^{\tilde{k}}]=\mathbf{R}(\mathcal{L}^{\tilde{k}})\times_{1}[X_{1}^{\tilde{k}}]^{\top}\times_{2}[X_{2}^{\tilde{k}}]^{\top}\times_{3}[X_{3}^{\tilde{k}}]^{\top}$. Since we have
$\|[\mathcal{O}^{\tilde{k}}]-[\tilde{\mathcal{O}}^{\tilde{k}}]\|_{F}$ $\leq\sqrt{c_{2}}\|[X_{1}^{\tilde{k}}]\|_{F}\|[X_{2}^{\tilde{k}}]\|_{F}\|[X_{3}^{\tilde{k}}]\|_{F}\|\mathcal{L}^{\tilde{k}}-\mathcal{L}^{\tilde{k}-1}\|_{F}$
@@
$\Phi(\mathcal{Z}^{k})-\Phi(\mathcal{Z}^{k+1})\geq c_{3}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}$
@@
$\sum^{\infty}_{k=0}\sum_{i=1}^{3}\left\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\right\|_{F}^{2}\leq\frac{1}{c_{3}}\Phi(\mathcal{Z}^{0}).$
@@
(22) $\lim_{k\to\infty}\|([\mathbf{I}_{m_{i}}]-[{X}_{i}^{k}]{[{X}_{i}^{k}]}^{\top})[H_{i}^{k}]\|=0.$

### v3

@@ 4 Convergence Analysis of The P-BCD Algorithm
Let $h:\mathbb{R}^{m\times n}\to\mathbb{R}$ be defined by $h(\boldsymbol{X})=\frac{1}{2}\|\boldsymbol{X}-\boldsymbol{P}\boldsymbol{Q}^{\top}\|_{F}^{2}$, where $\boldsymbol{Q}\in\mathbb{R}^{n\times m}$, $\boldsymbol{P}\in\mathbb{R}^{m\times m}$ and $m\geq n$. If $\boldsymbol{X}\in\mathbb{S}_{m,n}$ and
(18) $\bar{X}=\operatorname{Proj}_{\mathbb{S}_{m,n}}(\boldsymbol{X}-\tau(\boldsymbol{X}-\boldsymbol{P}\boldsymbol{Q}^{\top})),$
where $\tau\in(0,1)$, then we have $\bar{\boldsymbol{X}}\in\mathbb{S}_{m,n}$ and
(19) $h(\boldsymbol{X})-h(\bar{\boldsymbol{X}})\geq\frac{\tau^{-1}-1}{2(\tau^{-1}+1+\theta)^{2}}\|(\boldsymbol{I}_{m}-\boldsymbol{X}{\boldsymbol{X}}^{\top})\nabla h(\boldsymbol{X})\|_{F}^{2},$
where $\theta=\|\boldsymbol{P}\boldsymbol{Q}^{\top}\|_{2}$.
@@
Let $H$ be defined as in (7). Let $\{\mathcal{Z}^{k}\}$ be the sequence generated by Algorithm 1 and $[\boldsymbol{X}_{i}^{k+1}]$ be computed by (10). Then $[\boldsymbol{X}_{i}^{k+1}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and the following inequality holds:
(20) $\begin{split}&H([\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\sum_{i=1}^{3}\left\|([\boldsymbol{I}_{m_{i}}]-[\boldsymbol{X}_{i}^{k}]{[\boldsymbol{X}_{i}^{k}]}^{\top})[\boldsymbol{H}_{i}^{k}]\right\|_{F}^{2}\end{split}$
@@
(21) $\left\|[\boldsymbol{H}_{i}^{k}]^{\top}[\boldsymbol{X}_{i}^{k}]-{[\boldsymbol{X}_{i}^{k}]}^{\top}[H_{i}^{k}]\right\|_{F}\leq c_{4}\|[\boldsymbol{X}_{i+1}^{k}]-[\boldsymbol{X}_{i}^{k}]\|_{F},$
where $c_{3}=\frac{{\tilde{\alpha}_{\boldsymbol{X}}}^{-1}-1}{2({\tilde{\alpha}_{\boldsymbol{X}}}^{-1}+1+\theta_{\max})^{2}}>0$, $c_{4}=2(\alpha_{\boldsymbol{X}}\sqrt{n_{\max}}+\theta_{\max})>0$, $\theta_{\max}=\max_{ijk}$ $\|[\boldsymbol{P}_{i}^{k}]^{(j)}([\boldsymbol{Q}_{i}^{k}]^{(j)})^{\top}\|_{2}$, $n_{\max}=\max\{n_{1},n_{2},n_{3}\}$, and
$[\boldsymbol{H}_{1}^{k}]$ $=\nabla_{[\boldsymbol{X}_{1}]}H([\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$[\boldsymbol{H}_{2}^{k}]$ $=\nabla_{[\boldsymbol{X}_{2}]}H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$[\boldsymbol{H}_{3}^{k}]$ $=\nabla_{[\boldsymbol{X}_{3}]}H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
@@
$[\boldsymbol{H}_{i}^{k}]=([\boldsymbol{X}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}])[\boldsymbol{Q}_{i}^{k}]^{\top},\quad i=1,2,3.$
@@
It follows from Lemma 4.5 that $[\boldsymbol{X}_{i}^{k+1}]\in\mathbb{S}_{m_{i},n_{i}}$.
To show the first inequality holds, the update of $[\boldsymbol{X}_{i}^{k+1}]$ in (10) can be viewed as the iteration in (18) with $h([\boldsymbol{X}])=\frac{1}{2}\|[\boldsymbol{X}]-[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\|_{F}^{2}$, $[\boldsymbol{X}]=[\boldsymbol{X}_{i}^{k}]$, $[\bar{\boldsymbol{X}}]=[\boldsymbol{X}_{i}^{k+1}]$ and $\tau=\tilde{\alpha}_{\boldsymbol{X}}=\frac{\delta}{\delta+\alpha_{\boldsymbol{X}}}$. Then for $i=1$ we have
$H([\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$-H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})$
$=$ $\frac{1}{2}\|[\boldsymbol{X}_{i}^{k+1}][\boldsymbol{Q}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}]\|_{F}^{2}-\frac{1}{2}\|[\boldsymbol{X}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}]\|_{F}^{2}$
$=$ $\frac{1}{2}\|[\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\|_{F}^{2}-\frac{1}{2}\|[\boldsymbol{X}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\|_{F}^{2}$
$\geq$ $c_{ij}^{k}\left\|\left([\mathbf{I}_{m_{i}}]-[\boldsymbol{X}_{i}^{k}]{[\boldsymbol{X}_{i}^{k}]}^{\top}\right)\left([\boldsymbol{X}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)\right\|_{F}^{2},$
where $c_{ij}^{k}=\frac{{\tilde{\alpha}_{X}}^{-1}-1}{2({\tilde{\alpha}_{\boldsymbol{X}}}^{-1}+1+\|[\boldsymbol{P}_{i}^{k}]^{(j)}([\boldsymbol{Q}_{i}^{k}]^{(j)})^{\top}\|_{2})^{2}}\geq c_{3}>0$ and $\theta_{\max}$ is bounded, since the sequence $\{\mathcal{Z}^{k}\}$ is bounded. Using $[\boldsymbol{X}_{i}^{k}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and $[\boldsymbol{X}_{i}^{k+1}]\in[\mathbb{S}_{m_{i},n_{i}}]$, we can obtain the fourth line above and rewrite part of the last line as follows
$\left([\mathbf{I}_{m_{i}}]-[\boldsymbol{X}_{i}^{k}]{[\boldsymbol{X}_{i}^{k}]}^{\top}\right)\left([\boldsymbol{X}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)$
$=$ $\left([\mathbf{I}_{m_{i}}]-[\boldsymbol{X}_{i}^{k}]{[\boldsymbol{X}_{i}^{k}]}^{\top}\right)\left([\boldsymbol{X}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}]\right)[\boldsymbol{Q}_{i}^{k}]^{\top}$
$=$ $\left([\boldsymbol{I}_{m_{i}}]-[\boldsymbol{X}_{i}^{k}]{[\boldsymbol{X}_{i}^{k}]}^{\top}\right)[\boldsymbol{H}_{i}^{k}].$
@@
$\begin{split}&H([\boldsymbol{X}_{1}^{k}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\mathbf{I}_{m_{1}}]-[\boldsymbol{X}_{1}^{k}]{[\boldsymbol{X}_{1}^{k}]}^{\top}\right)[\boldsymbol{H}_{1}^{k}]\right\|_{F}^{2}.\end{split}$
@@
$\begin{split}&H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\boldsymbol{I}_{m_{2}}]-[\boldsymbol{X}_{2}^{k}]{[\boldsymbol{X}_{2}^{k}]}^{\top}\right)[\boldsymbol{H}_{2}^{k}]\right\|_{F}^{2}\end{split}$
@@
$\begin{split}&H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k}],[\mathcal{G}^{k}],\mathcal{L}^{k})-H([\boldsymbol{X}_{1}^{k+1}],[\boldsymbol{X}_{2}^{k+1}],[\boldsymbol{X}_{3}^{k+1}],[\mathcal{G}^{k}],\mathcal{L}^{k})\\ \geq&c_{3}\left\|\left([\boldsymbol{I}_{m_{3}}]-[\boldsymbol{X}_{3}^{k}]{[\boldsymbol{X}_{3}^{k}]}^{\top}\right)[\boldsymbol{H}_{3}^{k}]\right\|_{F}^{2}.\end{split}$
@@
Next, we show the second inequality holds. According to the update of $[\boldsymbol{X}_{i}^{k+1}]$ in (10), if we let
$[\boldsymbol{U}^{k}][\boldsymbol{\Sigma}^{k}][\boldsymbol{V}^{k}]^{\top}$ $=[\boldsymbol{X}_{i}^{k}]-\tilde{\alpha}_{\boldsymbol{X}}\left([\boldsymbol{X}_{i}^{k}]-[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)$
$=\frac{1}{\alpha_{\boldsymbol{X}}+\delta}\left(\alpha_{\boldsymbol{X}}[\boldsymbol{X}_{i}^{k}]+\delta[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right),$
then the update $[\boldsymbol{X}_{i}^{k+1}]$ $=[\boldsymbol{U}^{k}][\boldsymbol{V}^{k}]^{\top}$. Hence, we have
(22) $[\boldsymbol{X}_{i}^{k+1}]^{\top}\left(\alpha_{\boldsymbol{X}}[\boldsymbol{X}_{i}^{k}]+\delta[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)$ $=\left(\alpha_{\boldsymbol{X}}[\boldsymbol{X}_{i}^{k}]+\delta[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)^{\top}[\boldsymbol{X}_{i}^{k+1}]$
$=(\alpha_{\boldsymbol{X}}+\delta)[\boldsymbol{V}]^{k}[\boldsymbol{\Sigma}^{k}][\boldsymbol{V}^{k}]^{\top}.$
Then using $[\boldsymbol{X}_{i}^{k}]\in[\mathbb{S}_{m_{i},n_{i}}]$ and $[\boldsymbol{X}_{i}^{k+1}]\in[\mathbb{S}_{m_{i},n_{i}}]$, we can rewrite
$[\boldsymbol{H}_{i}^{k}]^{\top}[\boldsymbol{X}_{i}^{k}]-{[\boldsymbol{X}_{i}^{k}]}^{\top}[\boldsymbol{H}_{i}^{k}]$
$=$ $\left(\alpha_{\boldsymbol{X}}[\boldsymbol{X}_{i}^{k}]+\delta[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)^{\top}[\boldsymbol{X}_{i}^{k}]-[\boldsymbol{X}_{i}^{k}]^{\top}\left(\alpha_{\boldsymbol{X}}[\boldsymbol{X}_{i}^{k}]+\delta[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)$
$=$ $\left([\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{X}_{i}^{k}]\right)^{\top}\left(\alpha_{\boldsymbol{X}}[\boldsymbol{X}_{i}^{k}]+\delta[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)$
$-\left(\alpha_{\boldsymbol{X}}[\boldsymbol{X}_{i}^{k}]+\delta[\boldsymbol{P}_{i}^{k}][\boldsymbol{Q}_{i}^{k}]^{\top}\right)^{\top}\left([\boldsymbol{X}_{i}^{k+1}]-[\boldsymbol{X}_{i}^{k}]\right),$
where the last equation is obtained by (22). Taking the Frobenius norm of both sides, we obtain (21), where $2\|\alpha_{\boldsymbol{X}}[\boldsymbol{X}_{i}^{k}]^{(j)}+\delta[\boldsymbol{P}_{i}^{k}]^{(j)}([\boldsymbol{Q}_{i}^{k}]^{(j)})^{\top}\|_{F}\leq c_{4}$.
@@
${0}=\mathcal{L}^{\tilde{k}}+\mathcal{S}^{\tilde{k}}-\mathcal{D}+\delta\mathcal{W}_{\mathscr{R}}\odot\mathcal{L}^{\tilde{k}}-\delta\mathscr{R}^{\top}([\mathcal{Y}^{\tilde{k}}]),$
where $[\tilde{\mathcal{O}}^{\tilde{k}}]=\mathscr{R}(\mathcal{L}^{\tilde{k}})\times_{1}[\boldsymbol{X}_{1}^{\tilde{k}}]^{\top}\times_{2}[\boldsymbol{X}_{2}^{\tilde{k}}]^{\top}\times_{3}[\boldsymbol{X}_{3}^{\tilde{k}}]^{\top}$. Since we have
$\|[\mathcal{O}^{\tilde{k}}]-[\tilde{\mathcal{O}}^{\tilde{k}}]\|_{F}$ $\leq\sqrt{c_{2}}\|[\boldsymbol{X}_{1}^{\tilde{k}}]\|_{F}\|[\boldsymbol{X}_{2}^{\tilde{k}}]\|_{F}\|[\boldsymbol{X}_{3}^{\tilde{k}}]\|_{F}\|\mathcal{L}^{\tilde{k}}-\mathcal{L}^{\tilde{k}-1}\|_{F}$
@@
$\Phi(\mathcal{Z}^{k})-\Phi(\mathcal{Z}^{k+1})\geq c_{3}\sum_{i=1}^{3}\left\|([\boldsymbol{I}_{m_{i}}]-[\boldsymbol{X}_{i}^{k}]{[\boldsymbol{X}_{i}^{k}]}^{\top})[\boldsymbol{H}_{i}^{k}]\right\|_{F}^{2}$
@@
$\sum^{\infty}_{k=0}\sum_{i=1}^{3}\left\|([\boldsymbol{I}_{m_{i}}]-[\boldsymbol{X}_{i}^{k}]{[\boldsymbol{X}_{i}^{k}]}^{\top})[\boldsymbol{H}_{i}^{k}]\right\|_{F}^{2}\leq\frac{1}{c_{3}}\Phi(\mathcal{Z}^{0}).$
@@
(23) $\lim_{k\to\infty}\|([\boldsymbol{I}_{m_{i}}]-[\boldsymbol{X}_{i}^{k}]{[\boldsymbol{X}_{i}^{k}]}^{\top})[\boldsymbol{H}_{i}^{k}]\|=0.$